Find the indicated value of the log function:
log 6(36) 'The 6 is 'lower case'

Answer by stanbon(75887)

Find the indicated value of the log function:
log 6(36)
----
Let x = log6(36)
---
base = 6
exponent = x
result = 36
------
Exponential Form:
6^x = 36
6^x = 6^2
x = 2
------
log6(36) = 2
===================
Cheers,
Stan H.
===================